I really hope that someone can help me with this! Thank you in advance!!
A homeowner is installing a pool in his backyard, He wants its length to be 4 ft longer than its width. Then he wants to surround it with a concrete walkway 3 feet wide. If he can only afford 300 square feet of concrete for the walkway what should the dimensions for the pool be ?

Answer by scott8148(6628)

L = W + 4

(L + 6)(W + 6) - LW = 300

6L + 6W + 36 = 300

L + W = 44

substituting ___ (W + 4) + W = 44 ___ 2W = 40 ___ W = 20

substituting ___ L = (20) + 4 = 24